# JUHAN ARU, PHILÉMON BORDEREAU

### v1

@@ Acknowledgements
Both authors are supported by Eccellenza grant 194648 of the Swiss National Science Foundation and are members of NCCR Swissmap.
@@ 1.1. Notations
  • $\mathbb{D}$ denotes the unit disk in $\mathbb{C}$, $\mathbb{H}$ the upper half plane, and for $p>0$, $\mathbb{A}_{p}$ denotes the annulus $\mathbb{A}_{p}:=\left\{z\in\mathbb{C}:e^{-p}<\absolutevalue{z}<1\right\}$. A domain is a strict open subset of $\mathbb{C}$. We denote $D\subset\mathbb{C}$ for this strict inclusion.
  • Given $x\neq y\in\partial\mathbb{D}$, a domain $D\subset\mathbb{D}$ is an (non-crossing) admissible domain if it is finitely connected (in particular, connected) with the connected components of $\partial D$ given by $\mathcal{B}\cup\bigcup_{i=1}^{n}\mathcal{B}_{i}$, $n\geq 0$, with $\mathcal{B}$ and each $\mathcal{B}_{i}$ given by a Jordan curve. We further ask that $x,y\in\mathcal{B}\subset\partial D$ and that $D$ agrees with $\mathbb{D}$ in a neighborhood of $x$ and $y$. We denote by $(xy)$ or $\mathcal{B}_{r}$ (resp. $(yx)$ or $\mathcal{B}_{l}$) the counterclockwise arc from $x$ to $y$ (resp. from $y$ to $x$) in $\mathcal{B}$ (see also Figure 1).
    On the other hand, given $x\in\partial\mathbb{D}$ and $y\in\mathbb{D}$, a (crossing) admissible domain is a finitely connected domain $D\subset\mathbb{D}$ with $\partial D=\mathcal{B}_{\text{in}}\cup\mathcal{B}_{\text{out}}\cup\bigcup_{i=1}^{n}\mathcal{B}_{i}$, $n\geq 0$, with each component of $\partial D$ given by a Jordan curve. We also require that $x\in\mathcal{B}_{\text{out}}$ with $\mathcal{B}_{\text{out}}$ and $\partial\mathbb{D}$ agreeing locally around $x$, and that $y\in\mathcal{B}_{\text{in}}$ is analytic, i.e. that $\partial D$ is smooth at $y$.
    The requirement that boundary components correspond to Jordan curves is to avoid punctured domains. Furthermore, note that every admissible domain is conformally equivalent to a circle domain of the unit disk, i.e. such that $\mathcal{B}=\partial\mathbb{D}$, with each $\mathcal{B}_{i}$ corresponding to a circle in $\partial\mathbb{D}$: in particular we may assume that all boundary components are smooth, because the $\mathrm{SLE}$ measures that we construct can be transported by conformal equivalences.
  • Similarly, a domain $D^{\prime}$ is called a test domain (relatively to $(D,x,y)$ an admissible domain, crossing or non-crossing) if $D^{\prime}\subset D$, $D^{\prime}$ is simply-connected with $x,y\in\partial D^{\prime}=:\mathcal{B}^{\prime}$ and $D^{\prime}$ agrees with $D$ in a neighborhood of $x$ and $y$. We denote by $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$) the part of the counterclockwise arc from $y$ to $x$ (resp. from $x$ to $y$) in $\mathcal{B}^{\prime}$ that lies in $D$.
@@
  • For $\eta:[0,t_{\eta}]\to D$ a curve in a domain $D$ with $0\leq t_{\eta}\leq+\infty$, we write $\eta_{t}:=\eta([0,t])$ for $t\in[0,t_{\eta}]$. We abuse notation and write $\eta\subset D$ if $\eta(0,t_{\eta})\subset D$, allowing for $\eta(0),\eta(t_{0})\in\partial D$, which we call a curve from $\eta(0)$ to $\eta(t_{\eta})$ in $D$ (notice that the direction of time in the parametrization of the curve is implicit here). If $t<t_{\eta}$, we write $\eta_{t}\subset D$ when $\eta(0,t]\subset D$. If $\eta(0)$ and $\eta(t_{\eta})$ belong to the same boundary component, it separates $D$ in two parts we call $D_{l}$ or $(D\setminus\eta)_{\text{left}}$ (resp. $D_{r}$ or $(D\setminus\eta)_{\text{right}}$) the left (resp. right) connected component of $D\setminus\eta$ relatively to the orientation of $\eta$.
@@
  • For $D$ a simply-connected domain and $x\neq y\in\partial D$, we denote by $g_{D}$ the harmonic extension of the boundary values $-\lambda:=-\sqrt{\pi/8}$ (resp- $+\lambda$) on $(yx)$ (resp. on $(xy)$).
@@
    for $f,g$ differentiable. For $\delta\subset D$ a smooth path or curve, we denote by $(\cdot,\cdot)_{\delta}$ the scalar product on $L^{2}(\delta)$. We will often abuse notation and write $(f,g)_{\delta}$ when $f$ is a distribution on $\delta$ tested against a smooth enough function $g$.
@@ 1.2. Poisson kernels and operators
  • The Green’s function in a domain $D$ at $z\neq w\in D$ is the density at $w$ of a standard Brownian motion started at $z$, killed upon exiting $D$. This is normalized so that $G_{\mathbb{D}}(0,z)=-(2\pi)^{-1}\log\absolutevalue{z}$ |. We stress that this is a choice of normalization that might differ to some papers in the field, where the Green’s function would be defined as $\tilde{G}_{D}(0,z)=-\log\absolutevalue{z}$ |. It is convenient for us here as it yields $G_{D}=(-\Delta)^{-1}$ distributionally, without extra multiplicative terms. Note that the choice for the multiplicative constant of $G_{D}$ affects the value of $\lambda$ (choosing $G_{D}(x,y)=-c\log\absolutevalue{x-y}+o(1)$ as |$\absolutevalue{x-y}\to 0$ corresponds to $\lambda=\pi\sqrt{c}/2$).
  • For $D$ a domain, a point $w\in\partial D$ is $\partial D$-analytic if there exists a neighborhood $N$ in $\mathbb{C}$ and $\varphi:N\to\mathbb{C}$ conformal with $\varphi(w)=0,\varphi(N\cap D)=\varphi(N)\cap\mathbb{H}$. In an admissible domain $(D,x,y)$ as above, $x$ and $y$ are $\partial D$-analytic. For $z\in D$ and $w$ $\partial D$-analytic, we let $H_{D}(z,w)$ denote the Poisson kernel, defined as the density at $w$ of the measure induced by the exit point of $D$ of a Brownian motion started at $z$, i.e. for all $V\in\partial D\cap N$,
    $\mathbb{P}\left[B_{T_{D}}^{z}\in V\right]=\int_{V}H_{D}(z,w^{\prime})\mathrm{d}w^{\prime},$
    well-defined as $w$ is $\partial D$-analytic. Under this normalization,
@@
  • For $z,w$ both $\partial D$-analytic, we let $H_{\partial D}(x,y)$ denote the boundary or excursion Poisson kernel, defined by
    $H_{\partial D}(z,w)=\partial_{\nu_{z}}H_{D}(z,w)=\partial_{\nu_{w}}H_{D}(z,w),$
    with $\nu_{z}$ the inward normal vector at $z$. We stress again that the choice of normalization for the Poisson kernel carries over to the boundary Poisson kernel.
@@
$H_{D}(x,y)$ =|$=\absolutevalue{\varphi^{\prime}(y)}H_{\varphi(D)}(\varphi(x),\varphi(y)),$ $x\in D,y\in\partial D$
$H_{\partial D}(x,y)$ =|$=\absolutevalue{\varphi^{\prime}(x)}\absolutevalue{\varphi^{\prime}(y)}H_{\partial\varphi(D)}(\varphi(x),\varphi(y)),$ $x,y\in\partial D$
with $x$ and $\varphi(x)$ analytic on $\partial D$ and $\partial\varphi(D)$ whenever $x\in\partial D$, same for $y$, so that the quantities are well-defined.
We now relate the Green’s function and the Poisson kernels to the mass of Brownian measures [LW04, Law08]. We can decompose the law $\mu(z,t)$ of a Brownian motion $B$ running at speed $2$ (this is to later match the normalization of the Green’s function) started at $z$, of time-duration $t$ as
$\mu(z,t)=\int_{\mathbb{C}}\mu(z,w,t)\mathrm{d}w,$
and identify $\mu(z,w,t)$ as $(4\pi t)^{-1}\exp(-\absolutevalue{z-w}^{2}/(4t))$ times the probability measure corresponding to a Brownian bridge of time-duration $t$ from $z$ to $w$. In a domain $D\subset\mathbb{C}$, the Brownian path measure is defined by restriction to paths that remain in $D$: as such, the Brownian measures verify a restriction property, and a form of conformal covariance that is inherited from the conformal invariance of Brownian motion. Green’s function then naturally appears as the mass of the time-integrated measure
$\mu_{D}(z,w)=\int_{0}^{+\infty}\mu_{D}(z,w,t)\mathrm{d}t.$
The measures $\mu_{D}(z,z)$ are infinite for $z\in D\subset\mathbb{C}$, with divergences coming from small loops (this corresponds to the fact that $G_{D}(z,w)\sim(2\pi)^{-1}\log\absolutevalue{z-w}$ | as $z\to w$). Furthermore, they are measures on rooted loops, and forgetting the root leads to the Brownian loop measures
$\mu_{D}^{\text{loop}}=\int_{D}\int_{0}^{+\infty}\frac{1}{2t^{2}}\mu_{D}^{\#}(z,z,t)\mathrm{d}t\mathrm{d}z,$
with $\mu_{D}^{\#}(z,z,t)$ corresponding to the law of a Brownian bridge from $z$ to $z$ of time duration $t$ on the event that it stays in $D$. The importance of these measures comes from the fact that they satisfy the restriction property, but also–less obviously–conformal invariance [LW04]. For $K_{1},K_{2}$ two compacts of $D$ such that $\mathrm{dist}(K_{1},K_{2})>0$, the mass
@@
of loops in $D$ that intersect both $K_{1}$ and $K_{2}$ is finite and also a conformal invariant: for $\varphi:D\to\varphi(D)$ a conformal equivalence, $m_{D}(K_{1},K_{2})=m_{\varphi(D)}(\varphi(K_{1}),\varphi(K_{2}))$. This quantity is also related to the (zeta-regularized) determinant of the Laplacian, see Proposition 2.1 in [Dub09].
The Poisson and boundary Poisson kernels also describe the masses of the interior-to-boundary and boundary-to-boundary (excursions) Brownian measures. If $x,y\in\partial D$ are $\partial D$-analytic, then one can define the boundary-to-boundary or excursion measure as
$\mu_{\partial D}(x,y):=\lim_{\varepsilon\downarrow 0}\frac{1}{2\varepsilon^{2}}\mu_{D}(x+\varepsilon n_{x},y+\varepsilon n_{y}),$
a measure on paths from $x$ to $y$ in $D$. It can be realized as a measure of mass $H_{\partial D}(x,y)$ describing a Brownian excursion from $x$ to $y$, see once again [LSW03]. If $D^{\prime}\subset D$ agrees with $D$ around $x$ and $y$, then the definition of $\mu_{\partial D}(x,y)$ as a limit of interior Brownian measures readily shows that $\mu_{\partial D}(x,y)$ restricted to paths $\eta\subset D^{\prime}$ coincides with $\mu_{\partial D^{\prime}}(x,y)$.
We can also see the Poisson kernel and Green’s function as differential operators. For instance, for $\mathcal{B}_{1},\mathcal{B}_{2}$ disjoint smooth curves in a domain $D$, define $H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}:L^{2}(\mathcal{B}_{1})\to L^{2}(\mathcal{B}_{2})$ as
$H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}f=\int_{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}(\cdot,x)f(x)\mathrm{d}x.$
$(G_{D})_{\mathcal{B}_{i}}^{\mathcal{B}_{j}}$ can also be seen as an operator in the same way, but in this case allowing for $i=j$:
$((G_{D})_{\mathcal{B}_{i}}^{\mathcal{B}_{j}}f)(y)=\int_{\mathcal{B}_{i}}G_{D}(y,x)f(x)\mathrm{d}x\quad\text{ for }f\in L^{2}(\mathcal{B}_{i}),y\in\mathcal{B}_{j}.$
Similarly, we define $(H_{\partial D}-H_{\partial D\setminus\mathcal{B}_{2}})^{\mathcal{B}_{1}}:L^{2}(\mathcal{B}_{1})\to L^{2}(\mathcal{B}_{1})$ as
$(H_{\partial D}-H_{\partial D\setminus\mathcal{B}_{2}})^{\mathcal{B}_{1}}f=\int_{\mathcal{B}_{1}}(H_{\partial D}-H_{\partial D\setminus\mathcal{B}_{2}})(\cdot,x)f(x)\mathrm{d}x.$
We stress that only the difference of the boundary Poisson kernels is integrable (as for instance $H_{D}(x,y)\asymp(x-y)^{-2}$). We will often drop the superscripts when the context makes it clear.
We first gather two useful identities.
Lemma 1 (Some identities for loop measures).
Let $\mathcal{B}_{1},\mathcal{B}_{2}$ be disjoint smooth simple curves in $D$ with $\mathcal{B}_{1}$ smooth, and $x\neq y\in\mathcal{B}_{1}$. It holds that
(3) $(G_{D}-G_{D\setminus\mathcal{B}_{2}})(x,y)=\int_{\mathcal{B}_{2}}H_{D\setminus\mathcal{B}_{2}}(x,\mathrm{d}z)\int_{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}(z,\mathrm{d}w)G_{D}(w,y),$
(4) $\int_{\mathcal{B}_{2}}H_{D\setminus\mathcal{B}_{2}}(x,\mathrm{d}z)H_{D\setminus\mathcal{B}_{1}}(z,x)=\int_{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})(x,\mathrm{d}z)G_{D\setminus\mathcal{B}_{2}}(z,x).$
We might sometimes write more succintly these equations as
@@
$H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}$ $=(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})^{\mathcal{B}_{1}}(G_{D})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}.$
@@
The left-hand side in the first equation corresponds to the mass of Brownian paths from $x$ to $y$ in $D$ that hit $\mathcal{B}_{2}$. One can decompose such a path at its first hitting time of $\mathcal{B}_{2}$, and at its following first hitting time of $\mathcal{B}_{1}$: the mass of the measure disintegrated in this way is exactly the right-hand side.
The second equation follows similarly by the Markov property and reversibility of the Brownian measure: the left-hand side counts Brownian loops attached at $\mathcal{B}_{1}$ that ‘enter’ immediately $D\setminus\mathcal{B}_{1}$, hit $\mathcal{B}_{2}$, and come back to their starting point in $D$, which can be decomposed as an excursion from $\mathcal{B}_{1}$ to $\mathcal{B}_{1}$ in $D\setminus\mathcal{B}_{1}$ hitting $\mathcal{B}_{2}$ and a path to the starting point in $D\setminus\mathcal{B}_{2}$. ∎
Finally, we need a Proposition from [Dub09] that will be relevant in simplifying computations of Gaussian integrations:
Proposition 2 (Proposition 2.2 in [Dub09]).
@@
$\exp(-m_{D}(K_{1},K_{2}))=\mathrm{det}_{F}(I_{\mathcal{B}_{1}}-H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}),$
where $\mathrm{det}_{F}$ stands for the Fredholm determinant.
$I_{\mathcal{B}_{1}}$ stands here for the identity operator on $L^{2}(\mathcal{B}_{1})$. We will leave the notation $1_{\mathcal{B}_{1}}$ for the constant function in $L^{2}(\mathcal{B}_{1})$ equal to one.

### v2

@@ Acknowledgements
Both authors are supported by Eccellenza grant 194648 of the Swiss National Science Foundation and are members of NCCR Swissmap. The authors thank Baptiste Cerclé and anonymous referees for helpful comments.
@@ 1.1. Notations
[…]
@@
  • We call a simple curve in a domain $D$ an injective and continuous map $\eta:[0,t_{\eta}]\to D$ for some $0<t_{\eta}\leq+\infty$, and write $\eta_{t}:=\eta([0,t])$ for $t\in[0,t_{\eta}]$. We in fact allow curves to have endpoints on $\partial D$, i.e. that $\eta(0),\eta(t_{\eta})\in\partial D$, write $\eta\subset D$ if $\eta((0,t_{\eta}))\subset D$, and alternatively call the trace of those curves crosscuts from $\eta(0)$ to $\eta(t_{\eta})$ in $D$. When unspecified, simple curves (that we sometimes also call Jordan curves) are understood modulo increasing time-reparametrisation. If $\eta(0)$ and $\eta(t_{\eta})$ belong to the same boundary component, it separates $D$ in two parts we call $D_{l}$ or $(D\setminus\eta)_{\text{left}}$ (resp. $D_{r}$ or $(D\setminus\eta)_{\text{right}}$) the left (resp. right) connected component of $D\setminus\eta$ relatively to the orientation of $\eta$.
  • We call a (non-crossing) admissible domain a triplet $(D,x,y)$ where $D\subset\mathbb{D}$ is a connected domain whose boundary $\partial D$ can be written $\mathcal{B}\cup\bigcup_{j=1}^{n}\mathcal{B}_{j}$, $n\geq 0$, with $\mathcal{B}$ and each $\mathcal{B}_{j}$ given by a closed Jordan curve, and $x,y\in\mathcal{B}\cap\partial\mathbb{D}$. We further require that $\mathcal{B}$ agrees with $\partial\mathbb{D}$ in a neighbourhood of $x$ and $y$. We denote by $(xy)$ or $\mathcal{B}_{r}$ (resp. $(yx)$ or $\mathcal{B}_{l}$) the counterclockwise arc from $x$ to $y$ (resp. from $y$ to $x$) in $\mathcal{B}$ (see Figure 1 for a sketch).
    On the other hand, we call a (crossing) admissible domain the data $(D,x,y)$ where $D\subset\mathbb{D}$ is a finitely connected domain $D\subset\mathbb{D}$ with $\partial D=\mathcal{B}_{\text{in}}\cup\mathcal{B}_{\text{out}}\cup\bigcup_{j=1}^{n}\mathcal{B}_{j}$, $n\geq 0$, with each set in the union corresponding to a closed Jordan curve, together with $x\in\partial\mathbb{D}$ and $y\in\mathbb{D}$. We also require that $x\in\mathcal{B}_{\text{out}}$ with $\mathcal{B}_{\text{out}}$ and $\partial\mathbb{D}$ agreeing locally around $x$, and that $y\in\mathcal{B}_{\text{in}}$ is $\partial D$-smooth, i.e. that there is a neighbourhood $N$ of $y$ in $\mathbb{D}$ and a smooth map $\varphi:N\to\mathbb{C}$ with $\varphi(y)=0$ and $\varphi(D\cap N)=\mathbb{H}\cap\varphi(N)$. Both terminologies are meant to highlight when $x$ and $y$ belong to the same or different boundary components.
    The requirement that boundary components correspond to Jordan curves is to avoid punctured domains. Furthermore, note that by Koebe’s theorem every finitely connected domain is conformally equivalent to a circle domain of the unit disk, i.e. a domain of the form $\mathbb{D}\setminus C$, $C$ a finite union of closed disjoint disks in $\mathbb{D}$: in particular we will often work with admissible domains having smooth boundary, because the $\mathrm{SLE}$ measures that we construct can be transported by conformal equivalences. Importantly, these conformal equivalences can be extended to the boundary under this regularity assumption.
  • Similarly, a domain $D^{\prime}$ is called a test domain (relatively to $(D,x,y)$ an admissible domain, crossing or non-crossing) if $^{\prime}\subset D, DD^{\prime}$ is simply-connected with $x,y\in\partial D^{\prime}=:\mathcal{B}^{\prime}$ and $D^{\prime}$ agrees with $D$ in a neighbourhood of $x$ and $y$. We denote by $\mathcal{B}_{l}^{\prime}$ (resp. $\mathcal{B}_{r}^{\prime}$) the part of the counterclockwise arc from $y$ to $x$ (resp. from $x$ to $y$) in $\mathcal{B}^{\prime}$ that lies in $D$ (see again Figure 1). We will typically consider smooth test domains, i.e. test domains $D^{\prime}$ for which all boundary points are $\partial D^{\prime}$-smooth.
@@
  • For $D$ a smooth simply-connected domain and distinct $x,y\in\partial D$, we denote by $g_{D}$ the harmonic extension of the boundary values $\lambda:=-\sqrt{\pi/8}$ (resp. $\lambda$) on $(yx)$ (resp. on $(xy)$).
@@
    for $f,g\in H^{1}(D)$, the Sobolev space defined as the closure of $(C^{\infty}(D),\norm{\cdot}_{\nabla(D)})$. For $\delta\subset D$ a continuous curve, we denote by $(\cdot,\cdot)_{\delta}$ the scalar product on $L^{2}(\delta)$, with respect to the arclength measure on $\delta$ induced by the Euclidean metric on $\mathbb{C}$, that we will write $|\absolutevalue{\mathrm{d}x}$. We will often abuse notation and write $(f,g)_{\delta}$ when $f$ is a distribution on $\delta$ tested against a function $g$ of sufficient regularity.
@@ 1.2. Poisson kernels and operators
  • The Green’s function $G_{D}$ in a domain $D$ at distinct $z,w\in D$ (with Dirichlet boundary conditions) is the density at $w$ of the law of a Brownian motion with speed $2$ started at $z$, killed upon exiting $D$. This is normalised so that $|G_{\mathbb{D}}(0,z)=-(2\pi)^{-1}\log\absolutevalue{z}$, consistently with [43, 8], and so that $G_{D}$ is the integral kernel of $(-\Delta_{D})^{-1}$ without extra multiplicative terms.
  • Let $z\in D$ and $x$ be $\partial D$-smooth (meaning that there exists a neighbourhood $N$ of $x$ in $D$ and a smooth map $\varphi:N\to\mathbb{C}$ with $\varphi(x)=0$ and $\varphi(D\cap N)=\mathbb{H}\cap\varphi(N)$). We denote by $H_{D}(z,x)$ the Poisson kernel, defined as the density (w.r.t. arc length) at $x$ of the measure induced by the exit point of $D$ of a Brownian motion started at $z$, i.e. for all $V\in\partial D\cap N$,
    $|,\mathbb{P}\left[B_{T_{D}}^{z}\in V\right]=\int_{V}H_{D}(z,x^{\prime})\absolutevalue{\mathrm{d}x^{\prime}}$,
    well-defined as $x$ is $\partial D$-smooth. Under this normalisation,
@@
    The Poisson kernel also corresponds to the normal derivative of the Green’s function at the boundary, i.e.
    $H_{D}(z,w)=\partial_{\nu_{w}}G_{D}(z,w),$
    with $\nu_{w}$ the inward normal vector at $w$.
  • For distinct $x,y$ both $\partial D$-smooth, we let $H_{\partial D}(x,y)$ denote the boundary or excursion Poisson kernel, defined by
    $H_{\partial D}(x,y)=\partial_{\nu_{x}}H_{D}(x,y)=\partial_{\nu_{y}}H_{D}(x,y),$
    The boundary Poisson kernel also describes the mass of measures on Brownian excursions from the boundary of planar domains [23, Section 3.3]: this will be important later on for probabilistic interpretations of operators defined using these kernels.
@@
[…]
with $x$ and $\varphi(x)$ smooth on $\partial D$ and $\partial\varphi(D)$ whenever $x\in\partial D$, same for $y$, so that the quantities are well-defined.
An important family of measures on unrooted loops in domains of the complex plane are the Brownian loop measures
$\mu_{D}^{\text{loop}}\coloneqq\int_{D}\int_{0}^{+\infty}\frac{1}{4\pi t^{2}}\mu_{D}^{\#}(z,z,t)\mathrm{d}t\mathrm{d}A(z),$
with $\mu_{D}^{\#}(z,z,t)$ corresponding to the law of a Brownian bridge from $z$ to $z$ of time duration $t$, on the event that it stays in $D$. These measures satisfy a restriction property as well as conformal invariance [23, Proposition 6]. They are infinite, but we will make use of the fact that for $K_{1},K_{2}$ two compacts of $D$ such that $\mathrm{dist}(K_{1},K_{2})>0$, the mass
@@
of loops in $D$ that intersect both $K_{1}$ and $K_{2}$ is finite and also a conformal invariant: for $\varphi:D\to\varphi(D)$ a conformal equivalence, $m_{D}(K_{1},K_{2})=m_{\varphi(D)}(\varphi(K_{1}),\varphi(K_{2}))$.
We can also see the Poisson kernel and Green’s function as integral operators. For instance, for $\mathcal{B}_{1},\mathcal{B}_{2}$ disjoint smooth crosscuts in a domain $D$ with $\mathrm{dist}(\mathcal{B}_{1},\mathcal{B}_{2})>0$, define the bounded operator $H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}:L^{2}(\mathcal{B}_{1})\to L^{2}(\mathcal{B}_{2})$ as
$H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}f\coloneqq\int_{\mathcal{B}_{1}}f(x)H_{D\setminus\mathcal{B}_{1}}(\cdot,\mathrm{d}x),$
where $H_{D\setminus\mathcal{B}_{1}}$ corresponds to (abusing notation) the Poisson kernel in the connected component of $D\setminus\mathcal{B}_{1}$ that contains $\mathcal{B}_{2}$. In the same way, we can define integral operators $(G_{D})_{\mathcal{B}_{j}}^{\mathcal{B}_{k}},\left\{j,k\right\}\subseteq\left\{1,2\right\}$ out of the Green’s function and the boundary Poisson kernel for $f\in L^{2}(\mathcal{B}_{j})$ and $y\in\mathcal{B}_{k}$:
$((G_{D})_{\mathcal{B}_{j}}^{\mathcal{B}_{k}}f)(y)$ $|,\displaystyle\coloneqq\int_{\mathcal{B}_{j}}G_{D}(y,x)f(x)\absolutevalue{\mathrm{d}x}$,
[…]
$|(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}f\coloneqq\int_{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})(\cdot,x)f(x)\absolutevalue{\mathrm{d}x}$
[…]
We first gather useful identities between these operators.
Lemma 1.1 (Some identities for loop measures).
Let $\mathcal{B}_{1},\mathcal{B}_{2}$ be disjoint smooth simple curves in $D$ and distinct $x,y\in\mathcal{B}_{1}$. It holds that
[…]
When $\mathcal{B}_{2}$ is merely assumed continuous, it holds altogether that
[…]
We might sometimes write more succinctly these equations as
@@
$H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}$ $=(G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}},$
$(G_{D}-G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}$ $=(G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(G_{D})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}.$
@@
The left-hand side in Equation (5) corresponds to the mass of Brownian paths from $x$ to $y$ in $D$ that hit $\mathcal{B}_{2}$. One can decompose such a path at its first hitting time of $\mathcal{B}_{2}$, and at its following first hitting time of $\mathcal{B}_{1}$: the mass of the measure disintegrated in this way is exactly the right-hand side.
Equation 6 follows similarly by the Markov property and reversibility of the Brownian measure: the left-hand side counts Brownian loops started at $y$ that ‘enter’ immediately $D\setminus\mathcal{B}_{1}$, hit $\mathcal{B}_{2}$, and come back to a point $x\in\mathcal{B}_{1}$, which can be decomposed as an excursion from $\mathcal{B}_{1}$ to $\mathcal{B}_{1}$ in $D\setminus\mathcal{B}_{1}$ hitting $\mathcal{B}_{2}$ and a path to the starting point in $D\setminus\mathcal{B}_{2}$.
Finally, Equation 7 follows from both considerations, without any assumption on $\mathcal{B}_{2}$ other than continuity: a Brownian path from $x$ to $y$ intersecting $\mathcal{B}_{2}$ can be decomposed into a Brownian path from $x$ to some $w\in\mathcal{B}_{1}$ in $D\setminus\mathcal{B}_{2}$, an excursion from $w$ to some $z\in\mathcal{B}_{1}$ in $D\setminus\mathcal{B}_{2}$ intersecting $\mathcal{B}_{2}$, and a Brownian path from $z$ to $y$ in $D$. ∎
Finally, we need a proposition from [12] that will be relevant in simplifying computations of Gaussian integrations:
Proposition 1.2 (Proposition 2.2 in [12]).
@@
$\exp(-m_{D}(\mathcal{B}_{1},\mathcal{B}_{2}))=\mathrm{det}_{F}(I_{\mathcal{B}_{1}}-H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}),$
where $\mathrm{det}_{F}$ stands for the Fredholm determinant. In fact, it is sufficient that $\mathcal{B}_{1}$ is smooth and $\mathcal{B}_{2}$ continuous, provided that we understand
$H_{D\setminus\mathcal{B}_{2}}^{\mathcal{B}_{1}}H_{D\setminus\mathcal{B}_{1}}^{\mathcal{B}_{2}}:=(G_{D\setminus\mathcal{B}_{2}})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}(H_{\partial(D\setminus\mathcal{B}_{1})}-H_{\partial(D\setminus(\mathcal{B}_{1}\cup\mathcal{B}_{2}))})_{\mathcal{B}_{1}}^{\mathcal{B}_{1}}.$
$I_{\mathcal{B}_{1}}$ stands here for the identity operator on $L^{2}(\mathcal{B}_{1})$. We keep the notation $1_{\mathcal{B}_{1}}$ for the constant function in $L^{2}(\mathcal{B}_{1})$ equal to one. In [12], the result is stated for $\mathcal{B}_{2}$ smooth, but we explain in Appendix B how the result follows when $\mathcal{B}_{2}$ is merely continuous.